40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 112     Entry time: Sun May 2 19:18:36 2010
Author: Frank 
Type: Misc 
Category: RefCav 
Subject: reference cavity moved to new chamber 

i finally moved the reference cavity to the new chamber and pump it since Friday. The legs are now insulated and it comes with a removable, thick insulation using the good CertiFoam25 insulation.The sensors for stabilization are now glued to small pieces of Kapton tap which is sticks directly to the chamber surface. I'm still using the four AD590 sensors as we have all the electronics available and ready to use. I modified the temperature sensor readout box a little bit to gain a little more SNR. The gain in the first stage (TIA) is now 29.4k (~9.2V output). The individual outputs are now connected to the DAQ. In addition the sum of all four is subtracted from a reference voltage (AD586 reference) so that the output is 0 at 35C with a range of +-5K (we need at least 4.5K for tuning one FSR, so to not be stuck at either ends i made it a bit larger ). This signal is also connected to the DAQ. An additional sensor (PT1000)  is now connected to the top plate of the stack, right below the reference cavity. Readout electronics is in preparation, as well as for the other platinum sensors we can put on the outside of the chamber (not done yet). I'm also planning on replacing the noisy power supplies by low-noise current sources to drive the heaters. The required heating power could be reduced a lot to ~10W total now, measured with a not fully sealed insulation and missing parts. So i expect even less when we are done with that.

ELOG V3.1.3-